Circular No. 7944

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 O6 M. Suzuki, Utsunomiya, Tochigi, Japan, reports the discovery of a

comet on publicly accessible images taken with the SWAN instrument on the SOHO spacecraft on July 25 and 27 UT. Prediscovery images were found on images taken as early as July 13. Visual confirmation has been reported by A. Hale, Cloudcroft, New Mexico, with a 0.20-m reflector (coma diameter 4′; magnitude measurement hampered by twilight, moon and clouds). CCD frames taken by J. Broughton, Reedy Creek, Australia, with a 0.25-m reflector also confirm the object’s cometary nature.

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 13  1^{h}56^{m}  −34^{o}  12  Suzuki
16  2 08  −33  11  _{”}
18  2 11  −31  11  _{”}
20  2 25  −29  10  _{”}
25  2 53  −25  9  _{”}
27  3 10  −22  9  _{”}
Aug. 1.451  3 53.73  −13 58′.7  9.5 Hale
1.468  3 53.92  −13 56′.2  _{”}
1.72946 3 56 52.^{s}38 −13 15 29′′.2 13.2 Broughton
1.73076 3 56 53.28 −13 15 17.0 13.1  _{”}
1.73191 3 56 54.03 −13 15 06.5 13.1  _{”}

SUPERNOVA 2002eg IN UGC 11486 E. Sanders, Dominion Astrophysical Observatory, Victoria, British

Columbia, Canada, reports the discovery of an apparent supernova in R band CCD frames taken with the Plaskett 1.82-m reflector on July 26.34 (R mag ∼ 17.5) and 28.30 UT. W. D. Li, University of California at Berkeley, reports the confirmation of SN 2002eg on LOTOSS (cf. IAUC 7906) KAIT CCD frames on July 24.3 UT (mag ∼ 19), 29.3 (mag 18.1), and Aug. 1.3 (mag 17.7). The object does not appear in KAIT images on July 18.3 (limiting mag 19.5). The following position is from KAIT images: α = 19^{h}49^{m}47^{s}.25, δ = +50^{o}41′53′′.6 which is 15′′.1 west and 7′′.2 north of the center of UGC 11486. UGC 11486 is also the host galaxy of SN 2002dv (cf. IAUC 7930).

2002 August 1 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother